4^(2x+7)=32^(x+3)

Answer by stanbon(75887)

4^(2x+7)=32^(x+3)
----
[2^2]^(2x+7) = [2^5]^(x+3)
----
2^(4x+14) = 2^(5x+15)
----
4x+14 = 5x+15
x = -1
=======
Cheers,
Stan H.